OpenStudy (anonymous):

Can Someone Help . . . . . Please? Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 6)(x − 6) and (x − 6)(x − 6)

OpenStudy (souvik):

\[(a+b)*(a-b)=a ^{2}-b ^{2}\] and \[(a-b)*(a-b)=(a-b)^{2}\]

OpenStudy (anonymous):

so the first one would be x^2-36 and the second would be (x^2-12x+36)^2

OpenStudy (souvik):

last one would be (x^-12x+36)......no squared..!

OpenStudy (anonymous):

ok thanks.!

OpenStudy (souvik):

welcome..:)